11. Acadia Logistics anticipates that it will need more distribution center space to accommodate what it believes will be a significant increase in demand for its final-mile services. Acadia could either lease public warehouse space to cover all levels of demand or construct its own distribution center to meet a specified level of demand, and then use public warehousing to cover the rest. The yearly cost of building and operating its own facility, including the amortized cost of construction, is $11.50 per square foot. The yearly cost of leasing public warehouse space is $19.00 per square foot. See below for the expected demand requirements a. The expected value of leasing public warehouse space as required by demand is $ response as a whole number.) . (Enter your b. The expected value of building a 180,000-square-foot distribution center and leasing public warehouse space as required if demand exceeds the need for 180,000 square feet of space is $ whole number.) (Enter your response as a c. The expected value of building a 380,000-square-foot distribution center and leasing public warehouse space as required if demand exceeds the need for 380,000 square feet of space is $ whole number.) (Enter your response as a d. The expected value of building a 580,000-square-foot distribution center and leasing public warehouse space as required if demand exceeds the need for 580,000 square feet of space is $ whole number.) . (Enter your response as a e. The expected value of building a 780,000-square-foot distribution is $ number.) (Enter your response as a whole f. Which of the above decisions provides the minimized expected value? the minimum expected value 3: Expected demand requirementsrequired if demand exceeds the need for 580,000 square feet of space is $ whole number.) e. The expected value of building a 780,000-square-foot distribution is $ number) f. Which of the above decisions provides the minimized expected value? (Enter your response as a (Enter your response as a whole the minimum expected value 3: Expected demand Requirements (in sq. ft) 180,000 380,000 580,000 780,000 Probability 0.45 0.25 0.25 0.05 (1) O Both options B and D have Option E has O Both options A and C have O Option D has O Option B has O Both options A and B have Option C has
Answer #1

Expected requirement = =180000*0.45+380000*0.25+580000*0.25+780000*0.05 = 360000 sq ft

a) Expected value of leasing public warehouse = Expected requirement * Cost of leasing per sq ft = 360000 * 19 = $6840000

b) Expected value of building a warehouse of 180000  sq ft = 180000 * 11.50 = $2070000

Leasing required = Expected requirement - Warehouse built = (360000 - 180000)

Expected value of leasing in this case = (360000 - 180000) * 19 = $3420000

Total = $207000 + $342000 = $5490000

For options C,d and F since warehouse build > expected requirement. We will not need to calculate leasing requirement.

c)Expected value of building a warehouse of 380000 sq ft = 380000*11.5 = $4370000

d) Expected value of building a warehouse of 580000  sq ft =580000*11.5 = $6670000

e) Expected value of building a warehouse of 780000  sq ft = 780000*11.5 = $8970000

f) Option C has minimum expected value.